OpenStudy (anonymous):

Do these improper integrals converge or not? If they do, which are their values?

OpenStudy (abb0t):

Can you give an example? There are tests which you can use to determine whether the integral diverges or converges.

OpenStudy (anonymous):

Here are the integrals: \[\int\limits_{2}^{5}\frac{ 1 }{ \sqrt{x-2} }dx\] \[\int\limits_{-\infty}^{\infty}x.e ^{\frac{ x^{-2} }{ 2 }}dx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\frac{ 1 }{ x-3 }dx\]

OpenStudy (anonymous):

\[\begin{align*}\int_2^5\frac{dx}{\sqrt{x-2}}&=\lim_{t\to2^+}\int_t^5(x-2)^{-1/2}~dx\\ &=\lim_{t\to2^+}\left[2\sqrt{x-2}\right]_t^5\\ &=2\sqrt{5-2}-\lim_{t\to2^+}(2\sqrt{t-2}) \end{align*}\]

OpenStudy (anonymous):

\[\int_{-\infty}^\infty x~e^{-\frac{x^2}{2}}~dx\] Let \(u=-\dfrac{x^2}{2}\), so that \(-du=x~dx\). Keep in mind the change in the limits: as \(x\to\infty\), you have \(u\to-\infty\), and as \(x\to-\infty\), you have \(u\to-\infty\). So you have \[\int_{-\infty}^{-\infty}e^u~(-du)\] What do you know about definite integrals with upper/lower limits that are the same?

OpenStudy (anonymous):

\[\begin{align*}\int_0^3\frac{dx}{x-3}&=\lim_{t\to3^-}\int_0^3\frac{dx}{x-3}\\ &=\lim_{t\to3^-}\left[\ln|x-3|\right]_0^t\\ &=\lim_{t\to3^-}\ln|t-3|-\ln3 \end{align*}\]

OpenStudy (anonymous):

@JGMach, do you follow?

OpenStudy (anonymous):

Sorry for the stupid question but, why is it required to "put" a limit in the first one you solved?

OpenStudy (anonymous):

The function is undefined at \(x=2\). The area under the curve from some interval near 2 appears to be approaching \(\infty\); we don't know for sure. Using the limit allows to find the area under the curve as our interval slowly approaches this asymptote.